OpenStudy (anonymous):

2x + 3y = 7 in slope intercept form?

OpenStudy (amistre64):

'solve' for y

OpenStudy (anonymous):

how would I do that?

OpenStudy (amistre64):

algebra of course ... get y all by itself

OpenStudy (amistre64):

you want it to look like: y = _______

OpenStudy (anonymous):

so y = 2x + 7 - 3 ?

OpenStudy (amistre64):

close ... 2x + 3y = 7 -2x -2x ; subtract 2x ------------- 3y = -2x + 7 /3 /3 ; divide off that 3 -------------- y = -2/3 x + 7/3

OpenStudy (anonymous):

ohhhhhh okay :) thank you!

OpenStudy (amistre64):

:) good luck